# Electronic structure and transport in materials with flat bands: 2D materials and quasicrystals

### v1

@@ 2 Flat bands due to complex crystal geometry without defects
In condensed matter physics, electronic localization is often associated with a short-range confinement effect. This is particularly the case when local static defects or edge effects produce localized states and thus flat bands. We will study examples of these flat bands and their consequences on electrical conduction in chapter 3. However, it is now well established that complex structures without defects can also give rise to confined states with very low energy dispersion, i.e., with large lifetimes. This is thus the case in quasicrystals, their periodic approximants, and complex structures close to quasicrystals from the point of view of atomic order at medium and/or long distances. Since the 1990’s, it has been experimentally proven that this particular crystalline order can produce intermetallic alloys, such as AlPdRe, AlPdMn and AlCuFe, which are almost insulators [40, 45, 41, 47]. A characteristic of this new electronic localization due to quasiperiodic order is that the confined states are localized on a large number of atoms (several hundreds or thousands) in real space. Twisted bilayer 2D materials constitute another example of such electronic confinement by the crystalline structure. It has indeed been shown, first in TBG [90, 91, 92, 93], and then in other twisted bilayer 2D materials, e.g., Refs. [151, 152, 153, 154, 155, 156], that for some angles the lowest-energy states have a very low energy dispersion and therefore a very low velocity. As for quasicrystals and related phases, these states are spread over a very large number of atoms in the moiré cell resulting from the rotation angle between the two layers. In 2018, it was experimentally proven that this electronic localization by the moiré pattern can induce strong electronic correlations and a superconducting state for some fillings of the four flat bands [102, 103].
In this chapter, we present our work on quasiperiodic tilings (Sec. 2.2) and twisted bilayer 2D materials (Secs. 2.3 and 2.4) with emphasis on the properties of these confined states by defect-free crystallographic structures. In order to understand the analysis which has led to these results, it is important to present first (Sec. 2.1) the “Small Velocity Regime” of electronic states localized on a large number of atoms.
@@ 2.1 The Small Velocity Regime (SVR)
The concept of SVR was developed in 2006 to understand the anomalous conductivity of quasicrystals [27, 28, 157, 61] and complex intermetallic alloys [158] (Sec. 2.2). More recently we have shown that this concept also applies well to the quantum diffusion of charge carriers in the moiré pattern of twisted bilayer graphene near the magic angles (Sec. 2.3) [29], and to the universal conductivity of the midgap states in monolayer and bilayer graphene (Sec. 3.3.2) [159, 160].
@@
can be computed in the basis of Bloch states [27] (see A.2.1). In (2), $\langle\cdots\rangle_{E}$ is the average on all states at energy $E$. The average square spreading is the sum of two terms:
@@
$V_{\mathrm{B}}$ is the Boltzmann velocity in the $x$ direction. The semi-classical theory is equivalent to taking into account only this Boltzmann term. The second term, $\Delta X^{2}_{\mathrm{NB}}(E,t)$, is a non-ballistic (non-Boltzmann) contribution. This term refers to a bounded contribution (over time) that is associated with the minimal size $L_{w}$ of a wave packet, as discussed in the introduction. Roughly speaking, for a system without defects, $L_{w}$ at energy $E$ is the quadratic time average of the non-Boltzmann square spreading,
@@
When the eigenstates of a periodic system are known (by the Density Functional Theory (DFT) method or Tight Binding (TB) method), it is possible to calculate $\Delta X_{\mathrm{B}}$ and $\Delta X_{\mathrm{NB}}$ at any energy and any time (see A.2.1). The ballistic (Boltzmann) contribution and non-Boltzmann contribution are related to the intraband and interband matrix elements of the velocity operator, respectively. Note that the importance of interband terms, sometimes called “quantum” terms as opposed to “classical” intraband terms, has been already discussed in topological flat-band materials [12, 13, 15, 16, 9, 17, 18]. Some examples for $\Delta X_{\mathrm{B}}(t)$ and $\Delta X_{\mathrm{NB}}(t)$ are shown in Fig. 2.
Figure 2: Examples of Boltzmann and non-Boltzmann contributions to the average square spreading, $\Delta X^{2}=\Delta X_{\mathrm{B}}^{2}+\Delta X_{\mathrm{NB}}^{2}$, versus times $t$, for various energies $E$, in different periodic systems we have studied. (a) Realistic $\alpha$-AlMnSi approximant of quasicrystals: (a.1) zoom of total DOS, (a.2) $\Delta X^{2}$ for three energies close to Fermi energy $E_{F}$ (see (a.1)) [27, 157]. (b) Realistic 1/1-AlCuFe approximant at $E_{F}$ [157]. (c) $\lambda$-AlMn complex intermetallic alloy [158] that is not an approximant, but that has local and medium orders similar to an AlMn quasicrystal. The DOS of $\alpha$-AlMnSi and $\lambda$-AlMn are shown Fig. 3. (d) Octogonal tiling approximants: (d.1) at the same energy $E$, for approximants $k=3$, 4, 5, with increasingly large cell size (see Fig. 4); (d.2) approximant $k=5$ for many energies. TB calculations are done for a pure hopping potential [60]. Note that similar results have been obtained for the Penrose tiling [61]. (e) (25,26) Twisted bilayer graphene ($\theta=1.297^{o}$) close to the magic angle for two energies in the flat bands (see Sec. 2.3.3) [29]. $\Delta X_{\mathrm{NB}}^{2}(E,t)$ is calculated from Eq. (71). For (a), (b) and (c) eigenstates are calculated by DFT (LMTO [161]), whereas for (d) and (e), by TB models (see Refs.).
@@
Thus, when condition (6) is satisfied, the semi-classical (Boltzmann) approach for transport is no more valid, and a new diffusion regime, the small velocity regime (SVR), is reached. This makes it possible to draw up a simple phase diagram of the charge carrier transport regime (Fig. 1) as a function of the scattering length, $L=V_{\mathrm{B}}\tau$, due to elastic or inelastic defects. For realistic values of scattering time, $\tau\approx 10^{-14}$ s or $\tau\approx 10^{-13}$ s, in quasicrystals and approximants [43], DFT calculations have shown that the SVR is reached for many energies. Some results for $\alpha$-AlMnSi, $/1$ 1-AlCuFe and $\lambda$-AlMn are presented in Refs. [27, 157, 158], respectively, and Figs. 2 and 3(c). For magic angle twisted bilayer graphene (TBG), SVR is reached for states with energies corresponding to the low-energy flat-band states [29] (Fig. 2 and Sec. 2.3.4). The numerical results show different dependence of $L_{w}$ with the time $t$. Depending on the system and the energy, $L_{w}$ may increase with $t$ or saturate at large $t$. In this last case, one can define a new length,
@@
It is important to remark that $L_{w}$ can even diverge at large time $t$, as it is the case for monolayer graphene when $E\rightarrow E_{D}$ [29], without the SVR condition (6) being satisfied. However, for structures with large cells, such as approximants of quasicrystals and magic-angle TBGs, the SVR is achieved by a combined effect of a small velocity $V_{\mathrm{B}}$ and a large length $L_{w}$, due to the confinement of electrons by the medium order of the geometric structure with life time confinement of the order of magnitude of the scattering time $\tau$.
Figure 3: Electronic properties of actual Al-based quasicrystal approximants containing transition metals. (a) DFT (LMTO [161]) Density of states (DOS) of $\alpha$-AlMnSi approximants (140 atoms/cell) [162, 163, 157] and $\lambda$-AlMn (540 atoms/cell) complex intermetallic alloy, which is not a quasicrystal approximant, but its local atomic order is similar to that of AlMn quasicrystal [164]. (b) Variation of DOS due to Mn atoms in a metallic medium simulating an Al matrix, for 1 Mn atom (Friedel virtual bound state [165, 166]) and the Mn icosahedron commonly found in AlMn and AlPdMn quasicrystals and their approximants, and an icosahedron consisting of 12 Mn icosahedrons (144 atoms), built according to an inflation rule characteristic of periodic structures [167]. The spiky peaks are the signature of confined states by the clusters at both length scales. Their width, $\delta E$, corresponds to a lifetime of the order of $\hbar/\delta E\approx 10^{-13}$ s. (c) DC conductivity $\sigma$ versus the inverse scattering time $1/\tau$ in log–log scale. For each phase, numerical calculations are performed from DFT (LMTO) wave functions [158]. At room temperature, realistic scattering time $\tau$ values are about $10^{-14}$ s or $10^{-13}$ s [43], which corresponds to situations where non-Boltzmann terms may dominate (right part of the curves) in the approximants 1/1-AlCuFe, $\alpha$-AlMnSi, and the complex intermetallic alloys $\lambda$-AlMn.
@@ 2.2 Quantum diffusion in quasicrystals and quasiperiodic tilings
The present review is devoted to the localization mechanism that induces flat-band states. Rather than describe the properties of all quasicrystals, we will focus on Al-based quasicrystals, which are good examples of intermetallic alloys in which the atomic structure induces very specific localization at medium and long distances. Of course, the notion of band does not exist in quasicrystals, which have no translation periodicity. However, it is possible to consider approximate periodic structures, called “approximants”, with local and medium-range atomic orders similar to that of quasicrystals. It has been shown experimentally that approximants have electronic properties similar to those of quasicrystals [41, 53]. In the approximants, the bands are well defined and the flat bands correspond to a particular confinement of charge carriers by the quasiperiodic structure found also in the quasicrystals themselves. Experimentally, there are 3D quasicrystals, such as the icosahedral AlPdMn, AlCuFe and AlPdRe phases, and quasicrystals that are quasiperiodic in 2D and periodic in the third direction, such as the decagonal AlCuCo, AlCoNi phases. The localization mechanisms that are at the origin of the flat band in these complex materials are not directly related to the spatial dimension and apply to both 3D and 2D quasicrystals. It therefore seems appropriate to briefly review them here.
Experimental investigations have indicated that the conduction properties of many stable Al-based quasicrystals (AlCuFe, AlPdMn, AlPdRe, …) are unusual and differ strongly from those of more simple intermetallic alloys [42, 40, 45, 44, 46, 41, 43, 47, 168]. In particular, their conductivity increases when the static defects density increases and when the temperature increases (inversed Mathiessen rule). It appears also that the medium-range order and the chemical order –over one or a few nanometers– have a decisive influence. In particular it has been shown that the role of transition metals in the electronic and magnetic properties of aluminum based quasicrystals and more generally aluminum based complex intermetallic alloys is essential to reinforce the effect of the quasiperiodic potential on electronic states [162, 163, 169, 170, 171, 172, 173, 167, 174, 175, 176, 177, 178, 179]. There is now strong evidence that these non-standard properties result from the breakdown of the semi-classical Bloch-Boltzmann theory due to the SVR (Sec. 2.1). This phenomenon is presented in the following section (Sec. 2.2.1) for realistic approximants and related complex intermetallic phases. In Sec. 2.2.2, it is shown that long-range quasiperiodic order induces also SVR in Penrose tilings and octagonal 2D tilings without defects.
@@ 2.2.1 Quantum diffusion in aluminum-based approximants and complex intermetallic phases
The DOS of periodic approximants of aluminium-based quasicrystals and related complex phases (Fig. 3(a)) is characterized by two essential aspects. On the one hand, the presence of a pseudo-gap of typically 0.5 eV near the Fermi level contributes strongly to the stability of these phases by a Hume-Rothery mechanism [180], i.e., by a Fermi sphere lying on the Bragg planes associated with intense peaks in the diffraction images (pseudo-Brillouin zone) [40, 179, 181]. On the other hand, the presence of almost flat bands at many energies suggests that the DOS is composed of many spiky peaks, as predicted by DFT calculations [162, 163, 171, 174, 182, 164] and confirmed by surface spectroscopy experiments on quasicrystals [183, 184]. It has also been shown that these spiky peaks are the signature of states confined in the local atomic order characteristic of medium-range quasiperiodicity, in particular by clusters of transition metal atoms [167, 164] (Fig. 3(b)).
Figure 3(c) compares the conductivity calculated in approximants and complex phases with simple phases that have a standard metallic behavior. From equation (8), it is clear that the Boltzmann (non-Boltzmann) term increases (decreases) when $\tau$ increases. The minimum of diffusivity (conductivity) is thus obtained when, [158]
@@
For a scattering time $\tau$ such that $\tau>\tau^{*}$, Boltzmann terms dominate and the diffusivity (conductivity) increases as $\tau$ increases. Since $\tau$ decreases with increasing defect concentration and temperature, the behavior is metallic like: $\sigma$ decreases when defects and/or temperature increase. By contrast, for $\tau<\tau^{*}$, the conductivity increases when defects and/or temperature increase, and the behavior is insulating like. From DFT calculations (Fig. 3) in realistic phases, $\tau^{*}$ is around a few $10^{-14}$ or $\sim 10^{-13}$ s. These scattering time values correspond to scattering time estimates in quasicrystals and approximants from transport measurements at low temperature (4 K) [40, 45, 43, 47]. Therefore when the temperature increases from low values, these complex phases are insulating like, as experimentally found. From equation (9), when the Boltzmann term is negligible, i.e., $\tau\ll\tau^{*}$, the conductivity follows the inverse Mathiessen rule found experimentally [45, 43]:

### v2

@@ 2 Flat bands due to complex crystal geometry without defects
In condensed matter physics, electronic localization is often associated with a short-range confinement effect. This is particularly the case when local static defects or edge effects produce localized states and thus flat bands. We will study examples of these flat bands and their consequences on electrical conduction in section 3. However, it is now well established that complex structures without defects can also give rise to confined states with very low energy dispersion, i.e., with large lifetimes. This is thus the case in quasicrystals, their periodic approximants, and complex structures close to quasicrystals from the point of view of atomic order at medium and/or long distances. Since the 1990’s, it has been experimentally proven that this particular crystalline order can produce intermetallic alloys, such as AlPdRe, AlPdMn and AlCuFe, which are almost insulators Poon92 , Berger93 , Berger94 , Delahaye15 . A characteristic of this new electronic localization due to quasiperiodic order is that the confined states are localized on a large number of atoms (several hundreds or thousands) in real space. Twisted bilayer 2D materials constitute another example of such electronic confinement by the crystalline structure. It has indeed been shown, first in TBG Trambly10 , Bistritzer10 , SuarezMorell10 , Bistritzer11 , and then in other twisted bilayer 2D materials, e.g., Refs. Pan18 , Venky20 , Zhan20 , Ghiotto21 , Conti23 , Guo25 , that for some angles the lowest-energy states have a very low energy dispersion and therefore a very low velocity. As for quasicrystals and related phases, these states are spread over a very large number of atoms in the moiré cell resulting from the rotation angle between the two layers. In 2018, it was experimentally proven that this electronic localization by the moiré pattern can induce strong electronic correlations and a superconducting state for some fillings of the four flat bands Cao18a , Cao18b .
In this section, we present our work on quasiperiodic tilings (Sec. 2.2) and twisted bilayer 2D materials (Secs. 2.3 and 2.4) with emphasis on the properties of these confined states by defect-free crystallographic structures. In order to understand the analysis which has led to these results, it is important to present first (Sec. 2.1) the “Small Velocity Regime” of electronic states localized on a large number of atoms.
@@ 2.1 The Small Velocity Regime (SVR)
The concept of SVR was developed in 2006 to understand the anomalous conductivity of quasicrystals Trambly06 , Mayou08 , Trambly08 , Trambly14_QC and complex intermetallic alloys Trambly14_CRAS (Sec. 2.2). More recently we have shown that this concept also applies well to the quantum diffusion of charge carriers in the moiré pattern of twisted bilayer graphene near the magic angles (Sec. 2.3) Trambly16 , and to the universal conductivity of the midgap states in monolayer and bilayer graphene (Sec. 3.3.2) Bouzerar21 , Jouda21 .
@@
can be computed in the basis of Bloch states Trambly06 (see A.2.1). In (2), $\langle\cdots\rangle_{E}$ is the average on all states at energy $E$. The average square spreading is the sum of two terms:
@@
$V_{\mathrm{B}}$ is the Boltzmann velocity in the $x$ direction. The semi-classical theory is equivalent to taking into account only this Boltzmann term. The second term, $\Delta X^{2}_{\mathrm{NB}}(E,t)$, is a non-ballistic (non-Boltzmann) contribution. This term can refer to a bounded contribution (over time) that is associated with the minimal size $L_{w}$ of a wave packet, as discussed in the introduction. Roughly speaking, for a system without defects, $L_{w}$ at energy $E$ is the quadratic time average of the non-Boltzmann square spreading,
@@
When the eigenstates of a periodic system are known (by the Density Functional Theory (DFT) method or Tight Binding (TB) method), it is possible to calculate $\Delta X_{\mathrm{B}}$ and $\Delta X_{\mathrm{NB}}$ at any energy and any time (see A.2.1). The ballistic (Boltzmann) contribution and non-Boltzmann contribution are related to the intraband and interband matrix elements of the velocity operator, respectively. Note that the importance of interband terms, sometimes called “quantum” terms as opposed to “classical” intraband terms, has been already discussed in topological flat-band materials Mitscherling22 , Regnault2022 , Calugaru22 , Thumin23 , DanieliAndreanovLeykamFlach2024 , Thumin25 , Jiang25 . Some examples for $\Delta X_{\mathrm{B}}(t)$ and $\Delta X_{\mathrm{NB}}(t)$ are shown in Fig. 2.
Figure 2: Examples of Boltzmann and non-Boltzmann contributions to the average square spreading, $\Delta X^{2}=\Delta X_{\mathrm{B}}^{2}+\Delta X_{\mathrm{NB}}^{2}$, versus times $t$, for various energies $E$, in different periodic systems we have studied. (a) Realistic $\alpha$-AlMnSi approximant of quasicrystals: (a.1) zoom of total DOS, (a.2) $\Delta X^{2}$ for three energies close to Fermi energy $E_{F}$ (see (a.1)) Trambly06 , Trambly08 . (b) Realistic 1/1-AlCuFe approximant at $E_{F}$ Trambly08 . (c) $\lambda$-AlMn complex intermetallic alloy Trambly14_CRAS that is not an approximant, but that has local and medium orders similar to an AlMn quasicrystal. The DOS of $\alpha$-AlMnSi and $\lambda$-AlMn are shown Fig. 3. (d) Octogonal tiling approximants: (d.1) at the same energy $E$, for approximants $k=3$, 4, 5, with increasingly large cell size (see Fig. 4); (d.2) approximant $k=5$ for many energies. TB calculations are done for a pure hopping potential Trambly11_QC . Note that similar results have been obtained for the Penrose tiling Trambly14_QC . (e) (25,26) Twisted bilayer graphene ($\theta=1.297^{o}$) close to the magic angle for two energies in the flat bands (see Sec. 2.3.3) Trambly16 . $\Delta X_{\mathrm{NB}}^{2}(E,t)$ is calculated from Eq. (75). For (a), (b) and (c) eigenstates are calculated by DFT (LMTO Andersen75 ), whereas for (d) and (e), by TB models (see Refs.).
@@
Thus, when condition (6) is satisfied, the semi-classical (Boltzmann) approach for transport is no more valid, and a new diffusion regime, the small velocity regime (SVR), is reached. This makes it possible to draw up a simple phase diagram of the charge carrier transport regime (Fig. 1) as a function of the scattering length, $L=V_{\mathrm{B}}\tau$, due to elastic or inelastic defects. For realistic values of scattering time, $\tau\approx 10^{-14}$ s or $\tau\approx 10^{-13}$ s, in quasicrystals and approximants Mayou93 , DFT calculations have shown that the SVR is reached for many energies. Some results for $\alpha$-AlMnSi, $1/1$-AlCuFe and $\lambda$-AlMn are presented in Refs. Trambly06 , Trambly08 , Trambly14_CRAS , respectively, and Figs. 2 and 3(c). For magic angle twisted bilayer graphene (TBG), SVR is reached for states with energies corresponding to the low-energy flat-band states Trambly16 (Fig. 2 and Sec. 2.3.4). The numerical results show different dependence of $L_{w}$ with the time $t$. Depending on the system and the energy, $L_{w}$ may increase with $t$ or saturate at large $t$. In this last case, one can define a new length,
@@
It is important to remark that $L_{w}$ can even diverge at large time $t$, as it is the case for monolayer graphene when $E\rightarrow E_{D}$ Trambly16 , without the SVR condition (6) being satisfied. However, for structures with large cells, such as approximants of quasicrystals and magic-angle TBGs, the SVR is achieved by a combined effect of a small velocity $V_{\mathrm{B}}$ and a large length $L_{w}$, due to the confinement of electrons by the medium order of the geometric structure with life time confinement of the order of magnitude of the scattering time $\tau$.
Figure 3: Electronic properties of actual Al-based quasicrystal approximants containing transition metals. (a) DFT (LMTO Andersen75 ) Density of states (DOS) of $\alpha$-AlMnSi approximants (140 atoms/cell) Fujiwara89 , Fujiwara93 , Trambly08 and $\lambda$-AlMn (540 atoms/cell) complex intermetallic alloy, which is not a quasicrystal approximant, but its local atomic order is similar to that of AlMn quasicrystal Trambly09 . (b) Variation of DOS due to Mn atoms in a metallic medium simulating an Al matrix, for 1 Mn atom (Friedel virtual bound state Friedel56 , Anderson61 ) and the Mn icosahedron commonly found in AlMn and AlPdMn quasicrystals and their approximants, and an icosahedron consisting of 12 Mn icosahedrons (144 atoms), built according to an inflation rule characteristic of periodic structures Trambly97 . The spiky peaks are the signature of confined states by the clusters at both length scales. Their width, $\delta E$, corresponds to a lifetime of the order of $\hbar/\delta E\approx 10^{-13}$ s. (c) DC conductivity $\sigma$ versus the inverse scattering time $1/\tau$ in log–log scale. For each phase, numerical calculations are performed from DFT (LMTO) wave functions Trambly14_CRAS . At room temperature, realistic scattering time $\tau$ values are about $10^{-14}$ s or $10^{-13}$ s Mayou93 , which corresponds to situations where non-Boltzmann terms may dominate (right part of the curves) in the approximants 1/1-AlCuFe, $\alpha$-AlMnSi, and the complex intermetallic alloys $\lambda$-AlMn.
@@ 2.2 Quantum diffusion in quasicrystals and quasiperiodic tilings
The present review is devoted to the localization mechanism that induces flat-band states. Rather than describe the properties of all quasicrystals, we will focus on Al-based quasicrystals, which are good examples of intermetallic alloys in which the atomic structure induces very specific localization at medium and long distances. Of course, the notion of band does not exist in quasicrystals, which have no translation periodicity. However, it is possible to consider approximate periodic structures, called “approximants”, with local and medium-range atomic orders similar to that of quasicrystals. It has been shown experimentally that approximants have electronic properties similar to those of quasicrystals Berger94 , Mayou94 . In the approximants, the bands are well defined and the flat bands correspond to a particular confinement of charge carriers by the quasiperiodic structure found also in the quasicrystals themselves. Experimentally, there are 3D quasicrystals, such as the icosahedral AlPdMn, AlCuFe and AlPdRe phases, and quasicrystals that are quasiperiodic in 2D and periodic in the third direction, such as the decagonal AlCuCo, AlCoNi phases. The localization mechanisms that are at the origin of the flat band in these complex materials are not directly related to the spatial dimension and apply to both 3D and 2D quasicrystals. It therefore seems appropriate to briefly review them here.
Experimental investigations have indicated that the conduction properties of many stable Al-based quasicrystals (AlCuFe, AlPdMn, AlPdRe, …) are unusual and differ strongly from those of more simple intermetallic alloys Klein91 , Poon92 , Berger93 , Pierce93 , Akiyama93 , Berger94 , Mayou93 , Delahaye15 , Macia23 . In particular, their conductivity increases when the static defects density increases and when the temperature increases (inversed Mathiessen rule). It appears also that the medium-range order and the chemical order –over one or a few nanometers– have a decisive influence. In particular it has been shown that the role of transition metals in the electronic and magnetic properties of aluminum based quasicrystals and more generally aluminum based complex intermetallic alloys is essential to reinforce the effect of the quasiperiodic potential on electronic states Fujiwara89 , Fujiwara93 , Trambly93 , Mayou93_QC_d , Krajvci95 , Trambly95 , Trambly95_Ester , Trambly97 , Roche97_revue , Simonet98 , Hippert99 , Trambly00 , Trambly03 , Trambly05 . There is now strong evidence that these non-standard properties result from the breakdown of the semi-classical Bloch-Boltzmann theory due to the SVR (Sec. 2.1). This phenomenon is presented in the following section (Sec. 2.2.1) for realistic approximants and related complex intermetallic phases. In Sec. 2.2.2, it is shown that long-range quasiperiodic order induces also SVR in Penrose tilings and octagonal 2D tilings without defects.
@@ 2.2.1 Quantum diffusion in aluminum-based approximants and complex intermetallic phases
The DOS of periodic approximants of aluminium-based quasicrystals and related complex phases (Fig. 3(a)) is characterized by two essential aspects. On the one hand, the presence of a pseudogap of typically 0.5 eV near the Fermi level contributes strongly to the stability of these phases by a Hume-Rothery mechanism Massalski78 , i.e., by a Fermi sphere lying on the Bragg planes associated with intense peaks in the diffraction images (pseudo-Brillouin zone) Poon92 , Trambly05 , Mizutani12 . On the other hand, the presence of almost flat bands at many energies suggests that the DOS is composed of many spiky peaks, as predicted by DFT calculations Fujiwara89 , Fujiwara93 , Krajvci95 , Roche97_revue , NguyenManh03 , Trambly09 and confirmed by surface spectroscopy experiments on quasicrystals Widmer09 , Mader10 . It has also been shown that these spiky peaks are the signature of states confined in the local atomic order characteristic of medium-range quasiperiodicity, in particular by clusters of transition metal atoms Trambly97 , Trambly09 (Fig. 3(b)).
Figure 3(c) compares the conductivity calculated in approximants and complex phases with simple phases that have a standard metallic behavior. From equation (8), it is clear that the Boltzmann (non-Boltzmann) term increases (decreases) when $\tau$ increases. The minimum of diffusivity (conductivity) is thus obtained when, Trambly14_CRAS
@@
For a scattering time $\tau$ such that $\tau>\tau^{*}$, Boltzmann terms dominate and the diffusivity (conductivity) increases as $\tau$ increases. Since $\tau$ decreases with increasing defect concentration and temperature, the behavior is metallic like: $\sigma$ decreases when defects and/or temperature increase. By contrast, for $\tau<\tau^{*}$, the conductivity increases when defects and/or temperature increase, and the behavior is insulating like. From DFT calculations (Fig. 3) in realistic phases, $\tau^{*}$ is around a few $10^{-14}$ or $\sim 10^{-13}$ s. These scattering time values correspond to scattering time estimates in quasicrystals and approximants from transport measurements at low temperature (4 K) Poon92 , Berger93 , Mayou93 , Delahaye15 . Therefore when the temperature increases from low values, these complex phases are insulating like, as experimentally found. From equation (9), when the Boltzmann term is negligible, i.e., $\tau\ll\tau^{*}$, the conductivity follows the inverse Mathiessen rule found experimentally Berger93 , Mayou93 :